One number is 3 less than another. If 7 times the smaller number minus 3 times the larger number is 7, find the two numbers.
Answer by checkley75(3666)

X=Y+3
7Y-3X=7
7Y-3(Y+3)=7
7Y-3Y-9=7
4Y=7+9
4Y=16
Y=16/4
Y=4 ANSWER.
X=4=3
X=7 ANSWER.
PROOF
7*4-3*7=7
28-21=7
7=7